# Importance of Shot Noise in the Search for an Isotropic Stochastic Gravitational-Wave Background with Next Generation Detectors

### v1

@@ I Introduction
Ten years after the first direct detection of the gravitational wave (GW) event GW150914 [1], the observation of individually resolvable GW events has become increasingly routine. By the end of the third observation (O3) run of the LIGO-Virgo-KAGRA (LVK) collaboration, over 100 GW events have been detected [2] and a significant increase in this number is expected by the end of the fourth observation run (O4) [3, 4]. However, due to the limited sensitivity of current-generation ground-based GW detectors, including Advanced LIGO detectors at Hanford and Livingston [5], Advanced Virgo [6, 7] and KAGRA [8], we have not yet observed the stochastic gravitational wave background (SGWB), which arises as an incoherent superposition of innumerable GW signals across the whole history of the universe.
@@
Figure 1: A schematic comparison between the theoretical analysis and the practical analysis of the CBC foreground energy density. The blue bubble represents the “analysis in theory”, in which an theoretical expression for $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$ is shown. This integral has a rigorous one-to-one correspondence between population level hyperparameters $\bm{\Lambda}$ and $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$. The red bubble shows the “analysis in practice”, where the commonly used practical approximation is shown, which replaces the integral with a discrete summation over NN simulated events, hence depending on the exact simulation. Additionally, this approximation only holds when $T_{\mathrm{obs}}\to\infty$ or $N\to\infty$, whereas real observations are typically limited to a duration of $\mathcal{O}(1)$ year. The potential fluctuation of the summation given different realizations of NN events must therefore be evaluated. At the bottom, we show the cross-correlation estimator $\widehat{C}(f)$, which measures the observed energy density spectrum of the SGWB. The estimator $\widehat{C}(f)$ and corresponding uncertainty $\widehat{\sigma}(f)$ are defined in the Appendix C, and we refer to Ref. [46, 47] for more details. Due to the factors we list below “Cross-correlation $\widehat{C}(f)$”, $\widehat{C}(f)$ and $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$ can potentially deviate from each other. Here STFT refers to the short-time Fourier transform.
However, we emphasize that many past studies (including our previous work Ref. [44]) treat the CBC foreground as a persistent, Gaussian and stationary background. As we will discuss in detail here, this assumption does not accurately reflect the reality and, consequently, these analyses may underestimate the measurement uncertainties.
@@
$\Omega_{\mathrm{CBC}}(f|\mathbf{\Lambda})=\frac{f}{\rho_{c}}\int_{0}^{z_{\mathrm{max}}}\mathrm{d}z\frac{R(z)}{(1+z)H(z)}\Bigg\langle\frac{\mathrm{d}E_{\mathrm{GW}}}{\mathrm{d}f_{r}}\Big|_{f_{r}=f(1+z)}\Bigg\rangle.$ (1)
In the above expression, $f$ is the detector-frame frequency, $\bm{\Lambda}$ denotes the hyperparameters that determine the population details of the CBC system, $\rho_{c}$ is the critical energy density needed to close the universe, $z$ is the redshift, $R(z)$ is the source-frame merger rate of the corresponding CBC system, $H(z)$ is the Hubble parameter at the redshift $z$, $\frac{\text{d}E_{\mathrm{GW}}}{\text{d}f_{r}}$ is the energy emitted by the CBC system in the source frame at the source-frame frequency $f_{r}=f(1+z)$, and $\langle$$\rangle$ denotes ensemble average. Notably, both $R(z)$ and $\left\langle\frac{\text{d}E_{\mathrm{GW}}}{\text{d}f_{r}}\right\rangle$ implicitly depend on the hyperparameters $\bm{\Lambda}$, and therefore for a given set of hyperparameters $\bm{\Lambda}$, $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$ is fully determined, establishing a clear mapping from hyperparameters $\bm{\Lambda}$ to the SGWB energy density spectrum. However, the inverse mapping–from $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$ back to $\bm{\Lambda}$–may not be uniquely defined due to potential degeneracies among the hyperparameters.
@@
where $T_{\mathrm{obs}}$ denotes the observation time, and $\tilde{h}^{i}_{+/\times}(f)$ denotes the frequency domain waveform of the $i$th CBC event for $\times$ polarization. This approximation has been extensively employed in literature [49, 50, 51, 52, 53, 54]. Recently, Ref. [55] for the first time rigorously proved that this summation will finally converge to the integral when $T_{\mathrm{obs}}\to\infty$ or equivalently $N\to\infty$, which legitimates us to approximate the integral by this summation. However, we highlight that this result only holds when $T_{\mathrm{obs}}$ is large enough [^1] In practice, only when the statistical fluctuation of the foreground itself is much weaker than the detector noise can one claim $T_{\mathrm{obs}}$ is large enough. Therefore, the threshold of the long enough observational time $T_{\mathrm{th}}$ is a detector-related quantity.. In practical scenarios, where $T_{\mathrm{obs}}$ typically spans $\mathcal{O}(1)$ yr, statistical fluctuations in $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$, as computed by Eq. (2), across different realizations of the same underlying population model must be accounted for. The fluctuation introduces an additional uncertainty of the search, distinct from and complementary to the detector noise.
@@
In this work we focus on the detection of a SGWB in the XG scenario, where $\gtrsim 99\%$ BBHs are expected to be individually resolvable [57, 58]. Consequently, it may no longer be necessary to define a BBH foreground $\Omega_{\mathrm{BBH}}(f|\bm{\Lambda})$, as individual detections alone are informative enough to provide us with sufficient information about the entire population. In contrast, only around $50\%$ of BNS events are expected to be resolvable [57, 58]. Hence, in this work we focus on BNS systems and study the two new sources of uncertainty highlighted above. We emphasize that the methodologies we proposed herein are also applicable to both BBH and neutron star-black hole (NSBH) systems. We also refer interested readers to a recent work by Giarda et al. [59], which studies the uncertainty of the BBH foreground and uses a machine learning based method to estimate the corresponding uncertainty rapidly.
@@ II.2 Simulation Results
With BNS lists generated, we utilize Eq. (2) to compute the corresponding $\Omega_{\mathrm{BNS}}(f)$ for each realization. We present the results in Fig. 2, where each gray curve represents the computed $\Omega_{\mathrm{BNS}}(f)$ of a particular realization. Notably, several curves significantly deviate from the bulk distribution, exhibiting considerably higher energy densities. We highlight the three most prominent realizations using red, green and blue curves and indicate the redshift values of the nearest events in these particular cases. These realizations demonstrate that the resulting $\Omega_{\mathrm{BNS}}(f)$ can be dominated by nearby, exceptionally loud events, which is also known as the shot noise. While shot noise has been extensively explored in the context of the SGWB anisotropy in many studies, for instance, Ref. [62, 63, 64, 65], its impact on the search for an isotropic SGWB has yet to be thoroughly investigated. Although Ref. [51] addresses shot noise effects on the broadband estimator of $\Omega_{\mathrm{CBC}}(f_{\mathrm{ref}})$ at a reference frequency, it does not consider its implications for the narrowband estimator of $\Omega_{\mathrm{CBC}}(f)$.
@@
$\mathscr{L}(\widehat{C}(f)|\bm{\Lambda})\propto\exp\left(-\frac{(\widehat{C}(f)-\Omega_{\mathrm{BNS}}^{\mathrm{Model}}(f|\bm{\Lambda}))^{2}}{2\sigma_{n}^{2}(f)}\right),$ (7)
where $\sigma_{n}(f)$ is the corresponding uncertainty associated with $\widehat{C}(f)$. However, we now have shown that this expression could be problematic due to three reasons: (1) It is not proper to simply take the ensemble average $\langle\Omega_{\mathrm{BNS}}(f)\rangle$ as the definition of the quantity $\Omega_{\mathrm{BNS}}^{\mathrm{Model}}(f|\bm{\Lambda})$; (2) The above expression only considers the detector noise $\sigma_{n}(f)$, but ignores the statistical fluctuation of the BNS foreground itself considering the finite observational time; [^2] For the current generation of gravitational-wave detectors, due to the limited sensitivity, the statistical fluctuation of the foreground is expected to be weaker than the detector noise, so it is still safe to ignore this term, and still assume a Gaussian likelihood. [^3] If one does not aim to infer the population hyperparameters $\bm{\Lambda}$ of the corresponding foreground, but simply regard the foreground as a special realization of an ideal SGWB, which is persistent, stationary and Gaussian, then one does not need to consider the statistical fluctuation.(3) The deviations we show in Fig. 3 are not Gaussian, and hence the Gaussian likelihood itself may not hold in general.
@@ III Suppressing the Shot Noise by Removing Nearby Events
In Fig. 5, we focus on three particular frequencies: 10 Hz, 100 Hz and 200 Hz. For each $z^{\star}$ value we aggregate the relative deviations at these frequencies across all 200 realizations and plot the resulting distribution. Additionally, we fit a Gaussian to each of these three frequencies for comparison. We find that the deviation is not sensitive to frequency and that the overall distribution can be well approximated by a Gaussian if the nearby events are removed. Since we only simulate two hundred one-year-long realizations, the resulting distributions look not exactly the same as Gaussian distributions, whereas as the number of realizations and $T_{\mathrm{obs}}$ increase, the deviations are expected to converge to a Gaussian. A brief justification is as follows: if the observation time $T_{\mathrm{obs}}\to\infty$, and the number of BNS events $N\to\infty$, then $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)\approx N\langle\Omega_{\mathrm{BNS}}^{z^{\star}}(f)\rangle_{\mathrm{one~event}}$, where $\langle\Omega_{\mathrm{BNS}}^{z^{\star}}(f)\rangle_{\mathrm{one~event}}$ denotes the expected value of the energy density from a single BNS event. In this regime and assuming a particular population, the observed $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$ solely depends on $N$, which follows a Poisson distribution, i.e. $N\sim\mathrm{Poisson}[\lambda]$, where $\lambda$ is the expected value of $N$ given $T_{\mathrm{obs}}$. For large $\lambda$, this Poisson distribution approaches a Gaussian, $N\sim\mathcal{N}(\lambda,\lambda)$, and so does the fluctuations in $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$.
@@
In the left panel of Fig. 6, we compare the subtraction residual and the statistical fluctuation of the foreground, where the red curves correspond to the mean of the subtraction residual across the 200 realizations for different $z^{\star}$ values, and the orange curves show the standard deviation of $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$, which is denoted by $\sigma_{\Omega_{\mathrm{BNS}}^{z^{\star}}}(f)$ in the later discussion. The orientation of the arrow shows how red/orange curves change as $z^{\star}$ increases. As $z^{\star}$ increases starting from zero, the mean subtraction residual increases rapidly and surpasses the standard deviation of $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$ when $z^{\star}=0.04$. After this point, the mean subtraction residual plateaus as $z^{\star}$ approaches 0.50. This suggests that the subtraction residual is also dominated by nearby loud events, and thus also manifests as a form of shot noise.
@@
In the right panel of Fig. 6, we show the subtraction residual of 200 realizations for $z^{\star}=0.04$ case, where each gray curve corresponds to a particular realization. The $y$-axis shows the relative deviation of the subtraction residual of each realization from the mean subtraction residual. Unsurprisingly, we observe that the subtraction residual is indeed dominated by the “shot-noise” just like the original foreground itself, which makes its statistical description challenging. The loudest subtraction residual is around 14 times louder than the mean subtraction residual foreground. This phenomenon can be intuitively understood: when the original signal is very loud, even small deviations of the estimated parameters relative to the true parameters could result in a loud residual.
@@
As in Fig. 6, the arrows in Fig. 7 indicate the trends of how corresponding quantities change with increasing $z^{\star}$. We notice that the unresolvable BNS foreground increases with $z^{\star}$, while the mean and the standard deviation of $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$ decrease. In the above figure, we particularly plot the standard deviation of $\Omega_{\mathrm{BNS}}(f)$ in dash-dotted line corresponding to $z^{\star}=0$ case. We observe that if we keep all BNS events in the data, then the fluctuation of the foreground itself can be even stronger than the detector noise below 100 Hz, hence the uncertainty of the search can be dominated by $\sigma_{\Omega^{z^{\star}=0}_{\mathrm{BNS}}}(f)\equiv\sigma_{\Omega_{\mathrm{BNS}}}(f)$ instead of $\widehat{\sigma}_{n}(f)$. However, as long as we specify a $z^{\star}>0$, the corresponding $\sigma_{\Omega_{\mathrm{BNS}}^{z^{\star}}}(f)$ drops down by at least one order of magnitude and stays well below $\widehat{\sigma}_{n}(f)$. In addition, the unresolvable BNS foreground is roughly two orders of magnitude below the standard deviation of $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$, indicating that it can be safely neglected. As a sanity check, we also plot the estimated notching residual plus the unresolvable BNS foreground in brown for the special case $z^{\star}=0.35$. Because to perform notching procedure for all two hundred realizations is infeasible from the calculation aspect, we randomly draw five realizations and compute the mean of the notching residual plus the unresolvable foreground for them as the brown curve shown in Fig. 7. Please see Sec. IV for more details and see the following paragraph for the rationale of choosing $z^{\star}=0.35$.
Figure 8: The change of $\sigma_{\Omega_{\mathrm{BNS}}^{z^{\star}}}(f)$ at $f=f^{\star}=13$ Hz as a function of $z^{\star}$. The standard deviation decreases rapidly for small $z^{\star}$, then gradually levels off as $z^{\star}$ approaches about 0.35. The inset highlights this transition, showing that the rate of decrease becomes minimal beyond this point.
@@
Figure 9: Left: Cross-correlation spectrum $\widehat{C}(f)$ in dashed blue, theoretical estimate of the BNS foreground $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$ in green, and the cross-correlation uncertainty after notching, $\sigma_{n}^{\mathrm{N}}(f)$ in purple. The gray curves show the absolute difference between $\widehat{C}(f)$ and $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$, and the brown curves represent the notching residual plus the unresolvable foreground. The statistical fluctuation of the foreground is shown in orange. Middle: Relative deviation between $\widehat{C}(f)$ and $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$. Right: Impact on the total uncertainty when incorporating both the statistical fluctuation of the foreground and the deviation between cross-correlation and the theoretical estimate.

### v2

@@ I Introduction
Ten years after the first direct detection of the gravitational wave (GW) event GW150914 [1], the observation of individually resolvable GW events has become increasingly routine. By the end of the third observation (O3) run of the LIGO-Virgo-KAGRA (LVK) collaboration, over 90 GW events have been detected [2] and a significant increase in this number is expected by the end of the fourth observation run (O4) [3, 4]. However, due to the limited sensitivity of current-generation ground-based GW detectors, including Advanced LIGO detectors at Hanford and Livingston [5], Advanced Virgo [6, 7] and KAGRA [8], we have not yet observed the stochastic gravitational wave background (SGWB), which arises as an incoherent superposition of innumerable GW signals across the whole history of the universe.
@@
Figure 1: A schematic comparison between the theoretical analysis and the practical analysis of the CBC foreground energy density. The blue bubble represents the “analysis in theory”, in which an theoretical expression for $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$ is shown. This integral has a rigorous one-to-one correspondence between population level hyperparameters $\bm{\Lambda}$ and $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$. The red bubble shows the “analysis in practice”, where the commonly used practical approximation is shown, which replaces the integral with a discrete summation over NN simulated events, hence depending on the exact simulation. Additionally, this approximation only holds when $T_{\mathrm{obs}}\to\infty$ or $N\to\infty$, whereas real observations are typically limited to a duration of $\mathcal{O}(1)$ year. The potential fluctuation of the summation given different realizations of $N$ events must therefore be evaluated. At the bottom, we show the cross-correlation estimator $\widehat{C}(f)$, which measures the observed energy density spectrum of the SGWB. The estimator $\widehat{C}(f)$ and corresponding uncertainty $\widehat{\sigma}(f)$ are defined in the Appendix C, and we refer to Ref. [46, 47] for more details. Due to the factors we list below “Cross-correlation $\widehat{C}(f)$”, $\widehat{C}(f)$ and $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$ can potentially deviate from each other. Here STFT refers to the short-time Fourier transform.
However, we emphasize that many past studies (including our previous work Ref. [44]) treat the CBC foreground as a persistent, Gaussian and stationary background. As we will discuss in detail here, this assumption does not accurately reflect the reality and, consequently, these analyses may underestimate the measurement uncertainties. Before proceeding with a detailed discussion, we emphasize that in this work we use the term foreground to refer to the superposition of GW signals from all CBC events, including those that are individually resolvable.
@@
$\Omega_{\mathrm{CBC}}(f|\mathbf{\Lambda})=\frac{f}{\rho_{c}}\int_{0}^{z_{\mathrm{max}}}\text{d}z\,\frac{R(z)}{(1+z)H(z)}\Bigg\langle\frac{\mathrm{d}E_{\mathrm{GW}}}{\mathrm{d}f_{r}}\Big|_{f_{r}=f(1+z)}\Bigg\rangle.$ (1)
In the above expression, $f$ is the detector-frame frequency, $\bm{\Lambda}$ denotes the hyperparameters that determine the population details of the CBC system, $\rho_{c}=3H_{0}^{2}c^{2}/8\pi G$ is the critical energy density needed to close the universe, $z$ is the redshift, $R(z)$ is the source-frame merger rate of the corresponding CBC system, $H(z)$ is the Hubble parameter at the redshift $z$, $\frac{\text{d}E_{\mathrm{GW}}}{\text{d}f_{r}}$ is the energy emitted by the CBC system in the source frame at the source-frame frequency $f_{r}=f(1+z)$, and $\langle$$\rangle$ denotes ensemble average. Notably, both $R(z)$ and $\left\langle\frac{\text{d}E_{\mathrm{GW}}}{\text{d}f_{r}}\right\rangle$ implicitly depend on the hyperparameters $\bm{\Lambda}$, and therefore for a given set of hyperparameters $\bm{\Lambda}$, $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$ is fully determined, establishing a clear mapping from hyperparameters $\bm{\Lambda}$ to the SGWB energy density spectrum. However, the inverse mapping–from $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$ back to $\bm{\Lambda}$–may not be uniquely defined due to potential degeneracies among the hyperparameters.
@@
where $T_{\mathrm{obs}}$ denotes the observation time, $\tilde{h}^{i}_{+/\times}(f)$ denotes the frequency domain waveform of the $i$th CBC event for $\times$ polarization, and $N$ denotes the total number of events in a realization drawn from the population over an observation time $T_{\mathrm{obs}}$, accounting for both detectable and undetectable events. This approximation has been extensively employed in literature [49, 50, 51, 52, 53, 54]. Recently, Ref. [55] for the first time rigorously proved that this summation will finally converge to the integral when $T_{\mathrm{obs}}\to\infty$ or equivalently $N\to\infty$, which legitimates us to approximate the integral by this summation. However, we highlight that this result only holds when $T_{\mathrm{obs}}$ is large enough [^1] In practice, only when the statistical fluctuation of the foreground itself is much weaker than the detector noise can one claim $T_{\mathrm{obs}}$ is large enough. Therefore, the threshold of the long enough observational time $T_{\mathrm{th}}$ is a detector-related quantity.. In practical scenarios, where $T_{\mathrm{obs}}$ typically spans $\mathcal{O}(1)$ yr, statistical fluctuations in $\Omega_{\mathrm{CBC}}(f|\bm{\Lambda})$, as computed by Eq. (2), across different realizations of the same underlying population model must be accounted for. The fluctuation introduces an additional uncertainty of the search, distinct from and complementary to the detector noise.
@@
In this work we focus on the detection of a SGWB in the XG scenario, where $\gtrsim 99\%$ BBHs are expected to be individually resolvable [57, 58]. Consequently, it may no longer be necessary to define a BBH foreground $\Omega_{\mathrm{BBH}}(f|\bm{\Lambda})$, as individual detections alone are informative enough to provide us with sufficient information about the entire population. In contrast, only around $50\%$ of BNS events are expected to be resolvable [57, 58]. Hence, in this work we focus on BNS systems and study the two new sources of uncertainty highlighted above: (1) the energy density spectrum estimated by Eq. (2) can differ from the result given by Eq. (1) due to the fact that we do not have access to infinite number of sources in a period of observation time, and (2) the measured energy density spectrum based on the cross-correlation method can differ from the result predicted by Eq. (2) due to the subtleties we list above. We emphasize that the methodologies we proposed herein are also applicable to both BBH and neutron star-black hole (NSBH) systems. We also refer interested readers to a recent work by Giarda et al. [59], which studies the uncertainty of the BBH foreground and uses a machine learning based method to estimate the corresponding uncertainty rapidly.
@@ II.1 BNS Population Model
We highlight that this work aims to demonstrate the two new sources of uncertainty under the assumption of a particular underlying BNS population. If the assumed population model is changed, the results presented here may differ accordingly. We refer interested readers to Ref. [62] for a detailed discussion of how uncertainties in the population properties themselves propagate into uncertainties in $\Omega_{\mathrm{gw}}(f)$.
@@ II.2 Simulation Results
With BNS lists generated, we utilize Eq. (2) to compute the corresponding $\Omega_{\mathrm{BNS}}(f)$ for each realization, where we uese the Bilby [63] package to inject frequency domain waveform assuming IMRPhenomXAS [64] waveform model, which includes only the dominant quadrupole mode. As pointed out by Ref. [62], incorporating higher-order modes in the waveform has a negligible effect on the resulting $\Omega_{\mathrm{gw}}(f)$. We therefore do not consider the effect of higher-order modes in this work, and leave a more detailed investigation involving multiple waveform models to future studies. We present the results in Fig. 2, where each gray curve represents the computed $\Omega_{\mathrm{BNS}}(f)$ of a particular realization. Notably, several curves significantly deviate from the bulk distribution, exhibiting considerably higher energy densities. We highlight the three most prominent realizations using red, green and blue curves and indicate the redshift values of the nearest events in these particular cases. These realizations demonstrate that the resulting $\Omega_{\mathrm{BNS}}(f)$ can be dominated by nearby, exceptionally loud events, which is also known as the shot noise. While shot noise has been extensively explored in the context of the SGWB anisotropy in many studies, for instance, Ref. [65, 66, 67, 68], its impact on the search for an isotropic SGWB has yet to be thoroughly investigated. Although Ref. [51] addresses shot noise effects on the broadband estimator of $\Omega_{\mathrm{CBC}}(f_{\mathrm{ref}})$ at a reference frequency, it does not consider its implications for the narrowband estimator of $\Omega_{\mathrm{CBC}}(f)$.
@@
$\mathscr{L}(\widehat{C}(f)|\bm{\Lambda})=\frac{1}{\sqrt{2\pi}\sigma_{n}(f)}\exp\left(-\frac{(\widehat{C}(f)-\Omega_{\mathrm{BNS}}^{\mathrm{Model}}(f|\bm{\Lambda}))^{2}}{2\sigma_{n}^{2}(f)}\right),$ (7)
where $\sigma_{n}(f)$ is the corresponding uncertainty associated with $\widehat{C}(f)$. However, we now have shown that this expression could be problematic due to three reasons: (1) It is not proper to simply take the ensemble average $\langle\Omega_{\mathrm{BNS}}(f)\rangle$ as the definition of the quantity $\Omega_{\mathrm{BNS}}^{\mathrm{Model}}(f|\bm{\Lambda})$; (2) The above expression only considers the detector noise $\sigma_{n}(f)$, but ignores the statistical fluctuation of the BNS foreground itself considering the finite observational time; [^2] For the current generation of gravitational-wave detectors, due to the limited sensitivity, the statistical fluctuation of the foreground is expected to be weaker than the detector noise, so it is still safe to ignore this term, and still assume a Gaussian likelihood. This claim can be justified by Fig. 7 and by noting that the noise PSD of current detectors is approximately an order of magnitude larger than that of a CE detector, which in turn leads to an approximately two-order-of-magnitude increase in $\sigma(f)$., [^3] If one does not aim to infer the population hyperparameters $\bm{\Lambda}$ of the corresponding foreground, but simply regard the foreground as a special realization of an ideal SGWB, which is persistent, stationary and Gaussian, then one does not need to consider the statistical fluctuation.(3) The deviations we show in Fig. 3 are not Gaussian, and hence the Gaussian likelihood itself may not hold in general.
@@ III Suppressing the Shot Noise by Removing Nearby Events
In Fig. 5, we focus on three particular frequencies: 10 Hz, 100 Hz and 200 Hz. For each $z^{\star}$ value we aggregate the relative deviations at these frequencies across all 200 realizations and plot the resulting distribution. Additionally, we fit a Gaussian to each of these three frequencies for comparison. We find that the deviation is not sensitive to frequency for $z^{\star}>0$ cases and that the overall distribution can be well approximated by a Gaussian if the nearby events are removed. Since we only simulate two hundred one-year-long realizations, the resulting distributions look not exactly the same as Gaussian distributions, whereas as the number of realizations and $T_{\mathrm{obs}}$ increase, the deviations are expected to converge to a Gaussian. A brief justification is as follows: if the observation time $T_{\mathrm{obs}}\to\infty$, and the number of BNS events $N\to\infty$, then $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)\approx N\langle\Omega_{\mathrm{BNS}}^{z^{\star}}(f)\rangle_{\mathrm{one~event}}$, where $\langle\Omega_{\mathrm{BNS}}^{z^{\star}}(f)\rangle_{\mathrm{one~event}}$ denotes the expected value of the energy density from a single BNS event. In this regime and assuming a particular population, the observed $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$ solely depends on $N$, which follows a Poisson distribution, i.e. $N\sim\mathrm{Poisson}[\lambda]$, where $\lambda$ is the expected value of $N$ given $T_{\mathrm{obs}}$. For large $\lambda$, this Poisson distribution approaches a Gaussian, $N\sim\mathcal{N}(\lambda,\lambda)$, and so does the fluctuations in $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$.
@@
In the left panel of Fig. 6, we compare the subtraction residual and the statistical fluctuation of the foreground, where the red curves correspond to the mean of the subtraction residual across the 200 realizations for different $z^{\star}$ values, and the orange curves show the standard deviation of $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$, which is denoted by $\sigma_{\Omega_{\mathrm{BNS}}^{z^{\star}}}(f)$ in the later discussion. The orientation of the arrow shows how red/orange curves change as $z^{\star}$ increases. As $z^{\star}$ increases starting from zero, the mean subtraction residual increases rapidly and surpasses the standard deviation of $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$ when $z^{\star}=0.04$. After this point, the mean subtraction residual plateaus as $z^{\star}$ approaches 0.50.
@@
In the right panel of Fig. 6, we show the subtraction residual of 200 realizations for $z^{\star}=0.04$ case, where each gray curve corresponds to a particular realization. The $y$-axis shows the relative deviation of the subtraction residual of each realization from the mean subtraction residual. Unsurprisingly, we observe that the subtraction residual is indeed dominated by the “shot-noise” just like the original foreground itself, which makes its statistical description challenging. The loudest subtraction residual is around 14 times louder than the mean subtraction residual foreground. This phenomenon can be intuitively understood: when the original signal is very loud, one can obtain smaller fractional deviations of the estimated parameters from their true values; however, even such small fractional error in parameters may produce a large residual for a loud signal.
@@
As in Fig. 6, the arrows in Fig. 7 indicate the trends of how corresponding quantities change with increasing $z^{\star}$. We notice that the unresolvable BNS foreground increases with $z^{\star}$, while the mean and the standard deviation of $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$ decrease. In the above figure, we particularly plot the standard deviation of $\Omega_{\mathrm{BNS}}(f)$ in dash-dotted line corresponding to $z^{\star}=0$ case. We observe that if we keep all BNS events in the data, then the fluctuation of the foreground itself can be even stronger than the detector noise below 50 Hz, hence the uncertainty of the search can be dominated by $\sigma_{\Omega^{z^{\star}=0}_{\mathrm{BNS}}}(f)\equiv\sigma_{\Omega_{\mathrm{BNS}}}(f)$ instead of $\widehat{\sigma}_{n}(f)$. However, as long as we specify a $z^{\star}>0$, the corresponding $\sigma_{\Omega_{\mathrm{BNS}}^{z^{\star}}}(f)$ drops down by at least one order of magnitude and stays well below $\widehat{\sigma}_{n}(f)$. In addition, the unresolvable BNS foreground is roughly two orders of magnitude below the standard deviation of $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$, indicating that it can be safely neglected. As a sanity check, we also plot the estimated notching residual plus the unresolvable BNS foreground in brown for the special case $z^{\star}=0.35$. Because to perform notching procedure for all two hundred realizations is infeasible from the calculation aspect, we randomly draw five realizations and compute the mean of the notching residual plus the unresolvable foreground for them as the brown curve shown in Fig. 7. Please see Sec. IV for more details and see the following paragraph for the rationale of choosing $z^{\star}=0.35$.
Figure 8: The change of $\sigma_{\Omega_{\mathrm{BNS}}^{z^{\star}}}(f)$ at $f=f^{\star}=13$ Hz as a function of $z^{\star}$. The standard deviation decreases rapidly for small $z^{\star}$, then gradually levels off as $z^{\star}$ approaches about 0.35. The inset and the dashed oval highlight this transition, showing that the rate of decrease becomes minimal beyond this point.
@@
Figure 9: Left: Cross-correlation spectrum $\widehat{C}(f)$ in dashed blue, theoretical estimate of the BNS foreground $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$ in green, and the cross-correlation uncertainty after notching, $\sigma_{n}^{\mathrm{N}}(f)$ in purple. The gray curves show the absolute difference between $\widehat{C}(f)$ and $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$, and the brown curves represent the notching residual plus the unresolvable foreground. The statistical fluctuation of the foreground is shown in orange. Middle: Relative deviation between $\widehat{C}(f)$ and $\Omega_{\mathrm{BNS}}^{z^{\star}}(f)$. Right: Impact on the total uncertainty when incorporating both the statistical fluctuation of the foreground and the deviation between cross-correlation and the theoretical estimate. We fix $z^{\star}=0.35$ in this analysis.